# Test-Time Training Scaling Laws for Chemical Exploration in Drug Design

### v2

@@ A.3 MolRL-MGPT
The MolRL-MGPT model was re-implemented following the code provided from the original publication [36]. This CLM constituting the prior policy is a GPT-style Transformer model with 8 layers each with 8 attention heads. For a standardized comparison, the model was pre-trained on the benchmark training datasets respectively. For hardware reasons the batch size was lowered from 2048 to 512, with all other hyperparameters kept equal including training for 10 epochs. The learning rate, training loss, and SMILES validity during training is shown in x.
Appendix B Cooperative strategies
We employed the following cooperative strategies using for where the states and may be shared across agents in a cooperative manner. For brevity, we denote the probability of a full sequence generated as .
- Purge
- Shared
One shared replay buffer was used by all agents as a method of communicating high-rewarding areas of chemical space by on-policy augmentation. We anticipate this to result in convergence of agents to the same areas of chemical space and hence act more as a negative control.
- Shared with bonus
The same as Shared, however, an intrinsic reward was applied as a novelty bonus for molecules not already contained within the replay buffer. This reward shaping was applied before Equation 4.
(5) - Noise
Gaussian noise was applied to each agent’s parameters once at initialization. Such as to perturb each agent’s starting policy to encourage learning divergence. The effect of was tested on the validity of generated SMILES strings Figure 8. In cooperative experiments, we applied noise to all GRU hidden layers and the final linear layer.
(6) - RND
(7) The networks used in our implementation employ the same architecture as the CLM agents. The predictor network is trained to minimize the log-likelihood of SMILES sequences. The bonus is computed as the difference in policies between SMILES sequences measured by the summed log-likelihoods .
- ENTS
(8) The perplexities given the -th agent policy were summed over on-policy states of agents resulting in a state-conditioned policy uncertainty value for each set of agent states . The negative of this vector (so that smaller numbers represent more uncertainty) was softmax normalized resulting in a pseudo-probability of agent sampling states derived from different agents.
(9) - CES
(10) - DIFFS
An additional loss term for the -th agent to explicitly maximize the difference between uncertainty of states collected by previous agents relative to those collected by itself. Similar to ENTS the summed perplexity for collect states given the -th agent policy was calculated.
(11) - DIFFN
(12) - DvD
(13) where is the kernel value between agents and , is the squared Euclidean distance, and is a kernel bandwidth parameter.
(14) - POPNORM
Appendix C Molecular exploration benchmark
We selected four tasks, each with 2-4 objectives as molecular targets to be rediscovered based on real-world drug candidates ranging from pre-clinical (PC), clinical phases 1-3 (PH I-III) and approved (PH IV), shown in Figure 9 to Figure 12. For each task, the goal is to maximize the reward for a given molecule , where the reward is similarity to the closest molecular target :
@@ Appendix C Molecular exploration benchmark
Appendix D Baseline algorithms
@@
REINFORCE: This baseline was implemented as the vanilla reinforce algorithm where the objective is to maximize the likelihood of sequences weighted by the cumulative reward . In this implementation we additionally applied experience replay to augment on-policy data with previously high rewarded molecules. All configuration parameters are shown in Table 2.
REINVENT: This baseline was implemented following the REINVENT formulation [14] whereby a fixed version of the prior policy (i.e., the pre-trained CLM) is used to regularize the updates of the agent policy being trained. This is a form of reward shaping as shown in following Equation 20. This also introduces a hyperparameter that determines the balance between difference or the reward . The hyperparameters used corresponding to the default frequently used in the literature are shown in Table 3.
REINVENTMolOpt: This baseline is a hyperparameter optimized configuration of REINVENT for performance on the MolOpt benchmark. Notably this increases to focus more on molecule reward higher than regularization to the prior and increases the amount of experience replay used. All configuration parameter are shown in Table 4.
AHC: Augmented Hill-Climb [17] is a variation of REINVENT that conducts on-policy hill-climbing, i.e., on-policy molecules are ranked by reward and only the top- molecules are retained. This shift focuses learning on high-reward molecules which improves efficiency, and helps to avoid too much regularization in low-reward scenarios. All configuration parameters are shown in Table 5.
@@
Appendix E Baseline performance
E.1 MolExpL
E.2 MolExp
@@ E.3 GuacaMol
Appendix F TTT scaling
F.1 MolExpL

### v3

@@ A.3 MolRL-MGPT
The MolRL-MGPT model was re-implemented following the code provided from the original publication [37]. This CLM constituting the prior policy is a GPT-style Transformer model with 8 layers each with 8 attention heads. For a standardized comparison, the model was pre-trained on the benchmark training datasets respectively. For hardware reasons the batch size was lowered from 2048 to 512, with all other hyperparameters kept equal including training for 10 epochs. The learning rate, training loss, and SMILES validity during training is shown in Figure A.2.
Appendix B Molecular exploration benchmark
We selected four tasks, each with 2-4 objectives as molecular targets to be rediscovered based on real-world drug candidates ranging from pre-clinical (PC), clinical phases 1-3 (PH I-III) and approved (PH IV), shown in Figure B.1 to Figure B.4. For each task, the goal is to maximize the reward for a given molecule , where the reward is similarity to the closest molecular target :
@@ Appendix B Molecular exploration benchmark
Appendix C Baseline algorithms
A description of baseline algorithms and in particular their update step are detailed below, followed by the hyperparameters for each algorithm, followed by a pseudo-algorithm describing the general RL rollout and update procedure for these single-agent baseline algorithms 1. Note this algorithm does not apply to MolRL-MGPT which is a multi-agent baseline.
@@ Appendix C Baseline algorithms
REINFORCE: This baseline was implemented as the vanilla reinforce algorithm where the objective is to maximize the likelihood of sequences weighted by the cumulative reward . In this implementation we additionally applied experience replay to augment on-policy data with previously high rewarded molecules. All configuration parameters are shown in Table C.1.
REINVENT: This baseline was implemented following the REINVENT formulation [15] whereby a fixed version of the prior policy (i.e., the pre-trained CLM) is used to regularize the updates of the agent policy being trained. This is a form of reward shaping as shown in following Equation C.2. This also introduces a hyperparameter that determines the balance between difference or the reward . The hyperparameters used corresponding to the default frequently used in the literature are shown in Table C.2.
REINVENTMolOpt: This baseline is a hyperparameter optimized configuration of REINVENT for performance on the MolOpt benchmark. Notably this increases to focus more on molecule reward higher than regularization to the prior and increases the amount of experience replay used. All configuration parameter are shown in Table C.3.
AHC: Augmented Hill-Climb [18] is a variation of REINVENT that conducts on-policy hill-climbing, i.e., on-policy molecules are ranked by reward and only the top- molecules are retained. This shift focuses learning on high-reward molecules which improves efficiency, and helps to avoid too much regularization in low-reward scenarios. All configuration parameters are shown in Table C.4.
@@
Appendix D Baseline performance
D.1 MolExpL
D.2 MolExp
@@ D.3 GuacaMol
Appendix E TTT scaling
This section refers to test-time training scaling, in other words, multiple independent agents trained on the same task, which is generally described in Algorithm 2, further per-step details can be seen in 1.
E.1 MolExpL
@@ E.3 GuacaMol
Appendix F Cooperative strategies
We employed the following cooperative strategies using for where the states and may be shared across agents in a cooperative manner. For brevity, we denote the probability of a full sequence generated as . These cooperative strategies are adaptations to the general algorithm outlined in Algorithm 2.
- Purge
(F.1) - Shared
One shared replay buffer is used by all agents as a method of communicating high-rewarding areas of chemical space by on-policy augmentation. We anticipate this to result in convergence of agents to the same areas of chemical space and hence act as a negative control.
- Shared with bonus
The same as Shared, however, an additional intrinsic reward was applied as a novelty bonus for molecules not already contained within the replay buffer . This reward shaping was applied before other reward shaping mechanisms in Equation 2.5.
(F.2) - Noise
Gaussian noise was applied to each agent’s parameters once at initialization. Such as to perturb each agent’s starting policy to encourage learning divergence. The effect of was tested on the validity of generated SMILES strings Figure F.1. In cooperative experiments, we applied noise to all GRU hidden layers and the final linear layer.
(F.3) - RND
(F.4) The networks used in our implementation employ the same architecture as the CLM agents. The predictor network is trained to minimize the log-likelihood of SMILES sequences. The bonus is computed as the difference in policies between SMILES sequences measured by the summed log-likelihoods .
- ENTS
(F.5) The perplexities given the -th agent policy were summed over on-policy states of agents resulting in a state-conditioned policy uncertainty value for each set of agent states . The negative of this vector (so that smaller numbers represent more uncertainty) was softmax normalized resulting in a pseudo-probability of agent sampling states derived from different agents.
(F.6) - CES
(F.7) - DIFFS
An additional loss term for the -th agent to explicitly maximize the difference between uncertainty of states collected by previous agents relative to those collected by itself. Similar to ENTS the summed perplexity for collected states given the -th agent policy was calculated.
(F.8) - DIFFN
(F.9) - DvD
(F.10) where is the kernel value between agents and , is the squared Euclidean distance, and is a kernel bandwidth parameter.
(F.11) - POPNORM